OpenStudy (anonymous):

WILL GIVE MEDAL AND FOLLOW Which equation can be used to find the two numbers whose ratio is 4 to 3 and that have a sum of 42?

OpenStudy (anonymous):

@XxdemonchildXx

OpenStudy (anonymous):

@fallenangelorchid

OpenStudy (fallenangelorchid):

o-o

OpenStudy (uri):

Where are the equations?

OpenStudy (anonymous):

Um honestly you have the wrong person. I am not really good at math.

OpenStudy (anonymous):

4x – 3x = 42 (3/4)x=42 4x + 3x = 42 (4/3)x=42

OpenStudy (anonymous):

this one is the first one it got messed up : 4x – 3x = 42

OpenStudy (anonymous):

4x -3x = 42

OpenStudy (anonymous):

how do i figure this out?

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@cierra.whisper

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@Compassionate